The following information is for the inventory of mini-kettles at Concord Company Limited for the month of May:

Date Transaction Units
In
Unit
Cost
Total Units
Sold
Unit
Price
Total
May 1 Balance 140 $4.20 $588
6 Purchase 850 4.20 3,570
7 Sale 320 $7.20 $2,304
10 Sale 320 7.40 2,368
12 Purchase 420 5.00 2,100
15 Sale 240 8.00 1,920
18 Purchase 320 4.80 1,536
22 Sale 420 7.80 3,276
25 Purchase 500 4.40 2,200
30 Sale 240 7.60 1,824
Totals 2,230 $9,994 1,540

$11,692

A. Assuming that the periodic inventory method is used, calculate the inventory cost at May 31 under FIFO.

B. Assuming that the periodic inventory method is used, calculate the inventory cost at May 31 under weighted average. (Round weighted average unit cost to 2 decimal places, e.g. 52.75 and final answer to 0 decimal places, e.g. 5,275.)

C. Which method will yield the higher current ratio or gross profit?

______ will yield the highest ending inventory

______  gives a higher gross profit.
Answer #1

FIFO method states that goods purchased first are sold first

Weighted average method uses average cost for the purpose of calculation

Under periodic method, all records are updated at the end of the period and not after each transaction.

Units in Ending Inventory = 2230-1540 = 690 units

Inventory cost under FIFO = 500*4.4+190*4.8

= $3,112

Weighted average cost per unit = 9994/2230 = $4.48 per unit

Ending Inventory value = 4.48*690

= $3,091.2

FIFO will yield the highest ending inventory

And hence, highest current ratio

FIFO gives higher gross profit as cost of goods sold is lower